THE MICROWAVE ROTATIONAL SPECTRUM OF THE Ne-HCCCN VAN DER WAALS COMPLEX

\underline{AIKO HUCKAUF}, WOLFGANG JÄGER, and YUNJIE XU, Department of Chemistry, University of Alberta, Edmonton, AB T6G 2G2, Canada.

We report the results of the first rotational spectroscopic investigation of the Ne-HCCCN van der Waals complex. The measurements were carried out in the frequency range from 4 to 26 GHz using a molecular beam Fourier transform microwave spectrometer. HCCCN was prepared from ethyl propiolate in the usual way,^{a} and for our investigation we used a dilution of 0.2% ethynyl cyanide in neon at a total pressure of 3 to 4 atm. Strong b-type and weaker a-type transitions we observed. The spectra are consistent with a (vibrationally averaged) T-shaped geometry of the complex. The ^{14}N nuclear quadrupole hyperfine components of the rotational transitions were resolved. The spectral analyses yielded rotational, centrifugal distortion, and ^{14}N nuclear quadrupole coupling constants which were used to derive structural parameters and information about the large amplitude intermolecular motion.

aC. Moureau and J. C. Bongrand, Bull. Chem. Soc. 5, 846 (1909).